# WKB structure in a scalar model of flat bands

### v1

@@ 5.4. Mutliplicity in terms of the transition matrix
$\mathbf{u}=E_{+}v_{+}:=(\chi I_{+}+(1-\chi)I_{-})v_{+},\ \ \ u_{-}=E_{-+}v_{+}:=(I_{\mathbb{C}^{2}}-\mathcal{M}^{k}_{1\leftarrow 0})v_{+}.$ (5.61)
@@ 5.5. Quantisation condition
the eigenvalue equation can be written as
@@
Now, by Lemma 12
$\mathcal{M}_{\gamma(1)\leftarrow\gamma(0)}=B_{N}(\gamma(1);\alpha,h)^{-1}S_{\gamma}(\alpha,h)B_{N}(\gamma(0);\alpha,h)+\mathcal{O}_{\gamma,N}(h^{N+1}e^{c_{0}\operatorname{Im}\alpha/h}).$
@@
$B_{N}(\gamma(1);\alpha,h)=B_{N}(\gamma(0);\alpha,h),\quad S_{\gamma}(\alpha,h)=\begin{pmatrix}e^{-i\alpha h^{-1}c(h/\alpha)}&0\\ 0&e^{i\alpha h^{-1}c(h/\alpha)}\end{pmatrix}$.
@@
We can also assume that $|\operatorname{Re}k|\leq\pi h$ as conjugation by $e^{inx}$ means shifting $k$ by $nh$, $n\in\mathbb{Z}$.
@@
$F_{0}(\alpha,k,h):=\cos(c(\alpha/h)\alpha/h)-\cos(k/h)=0$ (5.65)
@@
We can now compare these solutions to the solutions of the equation (5.62), for $\operatorname{Im}\alpha\geq-h^{M}$, using (5.63) and Rouché’s theorem. The symmetry $\alpha\to-\alpha$ then gives the result for $\operatorname{Im}\alpha\leq 0$ as well. Also, the sign assumption is not needed as we can multiply $\alpha$ by a constant. Rescaling to eliminate $h$ we obtain the following theorem from which the rougher (no multiplicities) Theorem 3 follows:
@@
Under the assumptions of Theorem 3 there exist a contant $C_{0}$ and maps
@@
such that for $|\alpha|\geq C_{0}$, and in the notation of (5.6),

### v2

@@ 5.6. Mutliplicity in terms of the transition matrix
$\mathbf{u}=E_{+}v_{+}:=(\chi I_{+}+(1-\chi)I_{-})v_{+},\ \ \ u_{-}=E_{-+}v_{+}:=(\mathcal{M}^{k}_{0\leftarrow 1}-I_{\mathbb{C}^{2}})v_{+}.$ (5.62)
@@ 5.7. Quantisation condition
and using (5.13), the eigenvalue equation can be written as
@@
Now, by Lemma 12 (recalling (5.42))
$\mathcal{M}_{\gamma(1)\leftarrow\gamma(0)}=B_{N}(\gamma(1);\alpha,h)^{-1}S_{\gamma,N}(\alpha,h)B_{N}(\gamma(0);\alpha,h)+\mathcal{O}_{\gamma,N}(h^{N+1}e^{c_{0}\operatorname{Im}\alpha/h}).$
@@
$B_{N}(\gamma(1);\alpha,h)=B_{N}(\gamma(0);\alpha,h),\quad S_{\gamma,N}(\alpha,h)=\begin{pmatrix}e^{-i\alpha h^{-1}c(h/\alpha)}&0\\ 0&e^{i\alpha h^{-1}c(h/\alpha)}\end{pmatrix}$.
@@
We can also assume that $|\operatorname{Re}k|\leq\pi h$ since the spaces $\mathcal{V}_{k+2\pi h,h}$ and $\mathcal{V}_{k,h}$ defined in (5.47) are the same.
@@
$F_{0}(\alpha,k,h):=\cos(c(h/\alpha)\alpha/h)-\cos(k/h)=0$ (5.66)
@@
We can now compare these solutions to the solutions of the equation (5.63), for $\operatorname{Im}\alpha\geq-h^{M}$, using (5.64) and Rouché’s theorem applied to a contour consisting of $h^{N}$-sized disks centered at $\alpha_{\pm}(n,k,h)$. The symmetry $\alpha\to-\alpha$ then gives the result for $\operatorname{Im}\alpha\leq 0$ as well. To be consistent with this section, in comparison with Theorem 3 we assume that $\int_{\gamma}W(z)dz>0$, that is $\arg W_{0}=0$. Rescaling, $\alpha$ this can be removed.
@@
Under the assumptions of Theorem 3 and when $\arg W_{0}=0$, there exist a contant $C_{0}$ and maps
@@
and for $|\alpha|\geq C_{0}$, in the notation of (5.6),
@@
Acknowledgements. We would like to thank Zhongkai Tao for help with the material in §3.6, John Lott for the reference [Da07], and Zhenhao Li and Mengxuan Yang for comments on the first version of this paper.
The authors were partially supported by the Simons Targeted Grant Award No. 896630 “Moiré Materials Magic”, while the first named author was partially supported by NSF grant DMS-2400090 and a Simons Fellowship. Part of this work was completed while the first named author was in residence at the Mathematical Sciences Research Institute in Berkeley, California in Fall 2024, supported by NSF grant DMS-1928930.